OpenStudy (anonymous):

y = 4x and y + x = 5 (1 point) (–4, 1) (1, 4) (–3, 2) (2, 3)

jimthompson5910 (jim_thompson5910):

y + x = 5 4x + x = 5 ... replace y with 4x (since y = 4x) 5x = 5 x = 5/5 x = 1 now use this to find y (plug it into y = 4x)

OpenStudy (zpupster):

y+x=5 or y = 5-x 5-x = 4x or 5 = 5x x = 1

OpenStudy (anonymous):

its B?

jimthompson5910 (jim_thompson5910):

correct, if x = 1, then y = 4x y = 4*1 ... plug in x = 1 y = 4 So the solution is (1,4)